

********************************************************************
UARS WEEKLY STATUS REPORT - 15 MAR 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods.
***********
SUSIM
***********
On March 11th SUSIM was turned on after the previous yaw around maneuver.
Otherwise, the instrument performed its daily mid-resolution and
MgII solar scans during periods of predicted UARS telemetry.
We are preparing to make SUSIM generated attitude files available to
other UARS instruments. Stay tuned for details.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
A detailed review of the PEM CECIL PROCs has been completed. All 51 PEM
CECIL PROCs have been approved. PEM thanks Greg Haun for his assistance with
the CECIL conversion.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
PEM awaits support of the UARS Traceability Mission for the TIMED mission.
Initial meetings with a GUVI team representative have highlighted the need
for knowledge of conjunction periods. Investigation is underway into
identifying times when the PEM spectrometers and the GUVI view plane are
sampling from the same volume of the atmosphere. Orbital investigations
indicate that the region of time around April 28 looks promising. Initial
analysis of conjunction times indicate that the terminator is close to the
conjunction point. This effects the quality of the GUVI retrieval, which
works best in either total sunlight or total darkness. Additional
investigation is currently being conducted about this time period.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 02 March 2002 (Orbit 57259, GMT Day 061) through 08
March 2002 (Orbit 57363, GMT Day 067).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 063/2250 57303 Power OFF
PEM 065/2250 57333 Power OFF
SUSIM 065/2250 57333 Power OFF
HRDI 066/1553 57343 Power OFF
A Forward-to-Reverse Yaw Around maneuver (number 107) was successfully
executed in Orbit 57360 beginning 067/19:51:59Z and completing at
067/20:29:09Z on 08 March 2002. The SOLSTICE instrument was in day
only and all other instruments were in their off or standby mode
(including HRDI and PEM). Due to the SA position at 269 degrees, the
effective night length of the spacecraft after the maneuver was
approximately 57.1 minutes. Post maneuver analysis of data indicate
that during the Yaw Around, the end-of-night load bus voltage was
23.80 volts and the average depths of discharge (DODs) on Batteries 2
and 3 were approximately 21.44 and 21.20 percent respectively. The
next (reverse-to-forward) YAW is scheduled for 13 April 2002 (DOY
103).
The beta angle ranged between the angles of 25.33 degrees to 03.03
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 54 degrees on 27 March 2002 (DOY 086).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 24.48 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.94 to 1.31 degrees C. Additional
battery performance data for each battery on 02 March 2002 and 08
March 2002 are provided in Appendix D.
The clock error ranged from -11.6 to 7.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 6
HRDI = 2
HALOE = 1
SOLSTICE = 1
PEM = 1
SUSIM = 1
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 0
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/002 - 2287. 498845
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Eclipse testing remained the focal point this past week on UARS. The
integration of the new system has reached its final month and only a
few weeks are left before the April 1st deadline. The FOT continues
to test all of the Eclipse functionality against the UTTS and against
the spacecraft. Some errors have been found but the Raytheon software
development team has done an excellent job of turning around the fixes
for these problems. Thus the integration still remains on schedule.
A test from Eclipse to the spacecraft occurred on 05 March 2002. The
test performed was a partial table load without checksum and a full
table load of table 15. The test was successful and the table load
was verified via telemetry as well as via a dump and compare on the
Eclipse system.
MEETINGS
The FOT attended a UARS Eclipse Working Meeting on Thursday, March 07,
2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 13 April 2002 (DOY 103). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 08 March 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
57264 061/0832 02-066 HRDI TO DAY-ONLY OPS
57303 063/2250 02-067 HALOE Power OFF
57316 064/2031 02-068 OBC TBL 15 (ECLIPSE)
57332 065/2212 02-069 SOLSTICE FULL TIME TO DAY ONLY
57333 065/2250 02-070 PEM Power OFF
57333 065/2250 02-071 SUSIM Power OFF
57343 066/1553 02-072 HRDI Power OFF
57356 067/1232 02-073 OBC Clk Adj fr 0.0 to -20.6 msec/day
57360 067/1919 02-074 ACS TAM Removed fr Update (YAW)
57360 067/1919 02-075 ACS YAW FWD to REV
57361 067/2041 02-076 ACS TAM Alignment / Update Mode
57361 067/2041 02-075 ACS YAW GENASCII
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 08 March 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
_____________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 08 March 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
57356 067/1232 Clk Adj fr 0.0 to -20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 08 March 2002
02 March 2002 - GMT Day 061 Beta = 25.33 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -44.8/-61.6 -3.16/-4.38 +0.00/-0.00 10.72 0.000 0.00
2 +11.2/-89.6 +5.80/+5.14 +29.2/-13.6 25.12 1.030 21.20
3 +16.8/-56.0 +4.81/+3.82 +28.8/-13.6 25.12 1.030 20.86
08 March 2002 - GMT Day 067 Beta = 03.03 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 -39.2/-56.0 -3.77/-4.99 +0.0/ -0.0 10.72 0.000 0.00
2 +11.2/-95.2 +5.14/+4.48 +28.8/-13.2 24.48 1.030 20.97
3 +16.8/-67.2 +3.82/+3.17 +27.6/-12.8 24.48 1.010 20.86
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 08 March 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode